find the intercepts of the equation y=2x+2
Answer by ankor@dixie-net.com(22740)

find the intercepts of the equation y=2x+2
:
The y intercept is when x=0
y = 2(0) + 2
y = 2
:
The x intercept is when y = 0
2x + 2 = 0
2x = -2
x = -2/2
x = -1